OpenStudy (anonymous):

what is the solution to the rational equation (x/x^2-9)-(1/x+3)=(1/4x-12)

OpenStudy (kropot72):

\[\frac{x}{x ^{2}-9}-\frac{1}{x+3}=\frac{1}{4x-12}\] ???

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

You can factorise the denominator of the first term on the left hand side and factorise the denominator of the term on the right hand side.